this is a class i have to take before statistics, help me solve this question if two cards are drawn with out replacement from a deck, find the probability that the second is a spade, given that the the first card was a spade.
Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

The first card is a spade. So there are 13-1 = 12 spades left. On the second draw, there are 52 - 1 = 51 cards left.

So the probability of drawing a second spade given that the first is a spade is 12/51 = 4/17

Answer by stanbon(75887)

help me solve this question if two cards are drawn with out replacement from a deck, find the probability that the second is a spade, given that the the first card was a spade.
-----------
Once you have removed the 1st spade, there are 12 left.
P(draw spade after spade drawn) = 12/51
---------------------------------------
Cheers,
Stan H.